# references.bib

### v1

@@ Abstract
Learned visuomotor policies are capable of performing increasingly complex manipulation tasks. However, most of these policies are trained on data collected from limited robot positions and camera viewpoints. This leads to poor generalization to novel robot positions, which limits the use of these policies on mobile platforms, especially for precise tasks like pressing buttons or turning faucets. In this work, we formulate the policy mobilization problem: find a mobile robot base pose in a novel environment that is in distribution with respect to a manipulation policy trained on a limited set of camera viewpoints. Compared to retraining the policy itself to be more robust to unseen robot base pose initializations, policy mobilization decouples navigation from manipulation and thus does not require additional demonstrations. Crucially, this problem formulation complements existing efforts to improve manipulation policy robustness to novel viewpoints and remains compatible with them. To study policy mobilization, we introduce the Mobi-$\pi$ framework, which includes: (1) metrics that quantify the difficulty of mobilizing a given policy, (2) a suite of simulated mobile manipulation tasks based on RoboCasa to evaluate policy mobilization, (3) visualization tools for analysis, and (4) several baseline methods. We also propose a novel approach that bridges navigation and manipulation by optimizing the robot’s base pose to align with an in-distribution base pose for a learned policy. Our approach utilizes 3D Gaussian Splatting for novel view synthesis, a score function to evaluate pose suitability, and sampling-based optimization to identify optimal robot poses. We show that our approach outperforms baselines in both simulation and real-world environments, demonstrating its effectiveness for policy mobilization.
@@ 1 Introduction
Recent advances in learning visuomotor manipulation policies have enabled robots to execute increasingly complex manipulation tasks using imitation and reinforcement learning [chi2023diffusionpolicy, Zhao-RSS-23, brohan2022rt, brohan2023rt, team2024octo, kim2024openvla, black2024pi_0]. However, most manipulation policies are trained on data collected from a limited set of robot base poses and camera viewpoints, leading to a tight coupling between the policy’s input distribution and the configurations seen during training. When such a policy is deployed on a mobile robot, changes in the robot’s base pose inherently alter its viewpoint. This can cause both the visual input and the reachable workspace to drift outside the policy’s training distribution, severely degrading performance.
Given the abundance of stationary robot data, it is highly desirable to reuse existing datasets and pre-trained policies for mobile manipulation. While retraining a policy to handle a wide range of base poses and viewpoints is a straightforward solution, it is also costly—requiring extensive data collection across diverse positions and scene variations [brohan2022rt, khazatsky2024droid]. Moreover, as policies are deployed in mobile settings, their field of view expands, demanding stronger robustness to scene-level variations. An alternative strategy is to decouple navigation and manipulation by chaining a navigation module with the trained manipulation policy. However, most prior work in navigation specifies goals via language instructions or object locations [zhang2018semantic, rosinol2020kimera, chang2020semantic, ye2021auxiliary, deitke2022, gireesh2022object, gadre2022clip, ramrakhya2023pirlnav, hiroselelan, yokoyama2024vlfm] rather than selecting base poses that are policy-compatible. As a result, the robot often navigates to viewpoints that lie outside of the policy’s training distribution, leading to unreliable performance.
@@
To investigate the policy mobilization problem, we present the Mobi-$\pi$ framework. Mobi-$\pi$ provides tools for systematic study, including: (1) metrics to quantify the difficulty of mobilizing a given policy, (2) a suite of simulated mobile manipulation tasks built on RoboCasa [robocasa2024], (3), visualization tools for qualitative analysis, and (4) several baseline methods for benchmarking. We first evaluate non-policy aware baselines that decouple navigation and manipulation without considering that policies perform better from specific base poses. These approaches frequently fail due to the poor alignment with the policy’s training distribution. We also examine policy-aware baselines that learn to connect navigation and manipulation using large amounts of additional training data and show that they still struggle to generalize to unseen room layouts.
To address these limitations, we propose a method that searches for robot base poses likely to yield successful policy execution. Our approach (a) represents the scene using 3D Gaussian Splatting [kerbl3Dgaussians], (b) uses differentiable rendering to evaluate whether a candidate pose is in-distribution, provides visibility of task-relevant objects, and avoids collisions, and (c) optimizes over poses with a sampling-based procedure to identify the most suitable robot base location. We validate our approach on both the Mobi-$\pi$ simulation suite and three real-world mobile manipulation tasks. Our method outperforms the several baselines in various simulation and real-world experiments, enabling effective deployment of manipulation policies trained solely on stationary robot data.
@@ 2 Related Work
Semantic navigation. In semantic navigation, robots navigate environments to find specific objects or regions. Prior work [kuipers2000spatial, kuipers2004local, aydemir2013active, zhang2018semantic, zheng2019pixels, rosinol2020kimera, jin2023focusing] use classical methods like SLAM [kazerouni2022survey] to build and explore semantic maps, while recent methods use learning-based approaches [chang2020semantic, ye2021auxiliary, deitke2022, gireesh2022object, ramrakhya2023pirlnav, hiroselelan]. Other methods predict semantic top-down maps [ramakrishnan2022poni, chen2023not, yokoyama2024vlfm] and navigate using waypoint planners [hughes2022hydra]. These approaches typically stop in an area where the object is visible and close, without ensuring that the viewpoint is suitable for manipulation. In contrast, our method selects navigation targets that maximize downstream manipulation policy success.
@@
Imitation learning. Imitation learning teaches robots tasks from demonstrations [levine2016end, zeng2021transporter, shafiullah2022behavior, chi2023diffusionpolicy, aldaco2024aloha, fu2024mobile, chi2024universal], with recent efforts scaling to generalist models [brohan2022rt, brohan2023rt, gu2023rt, yang2024equivact, team2024octo, khazatsky2024droid, kim2024openvla, yang2024equibot, black2024pi_0]. Training such policies for both navigation and manipulation requires massive data. Instead, we use imitation learning for the manipulation policy and leverage existing navigation methods to move the robot to an in-distribution robot pose ensuring policy success. Our insight is that by selecting suitable poses for manipulation, we reduce the need for collecting more data from many robot poses or viewpoints and improve deployment in novel scenes. We focus on single-task policies and leave multi-task extensions to future work.
@@
Mobile manipulation. Prior work has studied various ways to bring learned policies to mobile manipulation [saycan2022arxiv, wu2023tidybot, shafiullah2023bringing, xiong2024adaptive, fu2024mobile, qiu2024wildlma, wu2024tidybot++, huang2024points2plans]. Some methods [brohan2022rt, fu2024mobile] rely on end-to-end learning but require large-scale data. Others [saycan2022arxiv, wu2023tidybot, shafiullah2023bringing, qiu2024wildlma] assume known initial base poses for a given task. ARPlace [2012_Freek] predicts valid robot poses for running skills but focuses more on reachability rather than viewpoint restrictions for these skills. ASC [yokoyama2023asc] trains coordinated mobile pick-and-place skills with RL in simulation but requires highly accurate simulators, which is often difficult to obtain for complex real-world tasks with deformables or complex multi-object interactions. Some works use reachability maps or abstract planners [jauhri2022robot, rosen2023synthesizing], but ignore that policies may be viewpoint sensitive. Our approach directly addresses this by finding robot base and camera poses that enable manipulation policy success. This “mobilizes” policies with fixed initial base poses, making them usable on mobile robots in complex environments, including those with deformable objects or multi-object interactions.
@@
In-distribution detection. Policy mobilization involves finding an in-distribution initial condition for a given policy. Prior work has extensively studied in-distribution detection or anomaly detection, covering classification-based methods [hendrycks2016baseline, liang2017enhancing, ruff2018deep, golan2018deep, fort2021exploring], reconstruction methods [hendrycks2018deep, nalisnick2018deep], methods that utilize generative adversarial networks (GANs) [di2019survey, schlegl2019f], and diffusion-based methods [liu2025survey]. Nearly all these methods assume a passive setting where they detect whether a given input is in-distribution without solving the problem of how an agent can find a sample that is more in-distribution. Recent robotics works have started to address out-of-distribution (OOD) detection [farid2022task, sinha2022system, agia2024unpacking] but treat the perception module as a passive component without actively updating the robot’s camera or base pose to mitigate OOD scenarios. Our work closes this gap, marrying in-distribution detection with active robot pose optimization in a mobile manipulation context.
@@
Hierarchical methods for policy learning. Our setting resembles Hierarchical Reinforcement Learning (HRL), where an option [sutton1999between] includes an initiation set, a policy, and termination function. HRL methods [bacon2017option, machado2017laplacian, frans2017meta, bagaria2021robustly, bagaria2023effectively] often learn these jointly but have limited success in complex settings. Instead of learning all elements of an option, we focus on selecting optimal robot poses for pre-trained manipulation policies in realistic environments.
@@
Training view-robust policies. Related to the problem setting of our work, prior work has studied methods of obtaining view-robust manipulation policies. One approach achieve this by collecting multi-view data in simulation [sadeghi2018sim2real, hirose2023exaug, seo2023multi], but suffers from the sim-to-real transfer challenge. Some works achieves view-robustness by using 3D representations, but training a policy to use 3D input requires data with well-calibrated camera extrinsics [gervet2023act3d, zhu2023learning, ze20243d], which is not readily available in many robot learning datasets today. Some works achieve real-world view robustness by manually collecting multi-view data in the real world [khazatsky2024droid], but this requires extensive human effort. There are proposals for using single-view novel view synthesis to bypass the need for multi-view data [tian2024viewinvariant], but methods for novel view synthesis from a single image struggle to effectively render views that are too different from the given viewpoint. From a view robustness perspective, our work takes advantage of the mobility of the robot so the robot can navigate to a base pose with an appropriate camera viewpoint for subsequent policy execution, thus requiring no extensive, additional data collection.
3 The Mobi-$\pi$ Framework
Given a visuomotor manipulation policy $\pi$ trained with imitation learning [chi2023diffusionpolicy, shafiullah2022behavior] on data recorded from one or a limited set of robot poses or viewpoints, we are interested in deploying $\pi$ on a mobile robot in an unseen scene $S$.
To study policy mobilization, we propose the Mobi-$\pi$ framework, which includes metrics to quantify the difficulty of mobilizing a given policy (Section 3.1), a suite of simulated mobile manipulation tasks to benchmark methods and policies for mobilization (Section 3.2), as well as baseline methods (Section 3.3).
3.1 Mobilization Feasibility Metrics
Given a policy for a task, we want to know how hard it is to mobilize it. We say a policy for a task is “difficult to mobilize” if running this policy on a mobile platform results in a large performance drop relative to its in-distribution performance. To quantify mobilization feasibility, we propose two distinct and complementary metrics as part of the Mobi-$\pi$-framework.
Visual mobilization feasibility metric. The visual appearance of the tasks can also affect mobilization difficulty. For example, if a policy manipulates an object while barely seeing the object in view, it will be harder to find a good location for policy execution. To quantify this, we evaluate how relevant objects of the task are represented in the scene. Formally, we sample $N_{v}$ images of the scene at near-optimal poses and compute the average percentage $S_{v}$ of the images occupied by the objects of interest in the task. The visual mobilization feasibility metric can be evaluated for methods that consider aligning the robot’s viewpoint with the policy’s capabilities during base placement. Naive baselines like LeLaN [hiroselelan] and VLFM [yokoyama2024vlfm] are not policy-aware, so this metric does not apply.
Figure 4: A suite of simulated tasks for benchmarking performance of policy mobilization methods. We pick five single-stage manipulation tasks from the RoboCasa benchmark [robocasa2024]: (1) Close Door: push or close the microwave or cabinet door using the robot gripper. (2) Turn on Faucet: precisely move the faucet handle to turn on water. (3) Turn on Microwave: push a small button on the microwave. (4) Close Drawer: close an open drawer next to the robot. (5) Turn on Stove: turn a specific knob out of six knobs on the stove. Here, we visualize two successful rollouts for each task. Note that episodes are executed in environments with different layouts and textures. In the Close Door task, the robot might encounter either a cabinet door or a microwave door. The Turn on Faucet, Turn on Microwave, and Turn on Stove tasks require accurate robot placement to be executed correctly. In the Close Drawer task, the drawer can be either on the left or right side of the robot.
3.2 Simulated Task Suite for Evaluating Policy Mobilization
Part of the Mobi-$\pi$ framework is a suite of simulated mobile manipulation tasks based on RoboCasa [robocasa2024] that will allow us to study the policy mobilization problem and benchmark different methods. Our task suite includes five challenging tasks illustrated in Figure 4. For each task, we train a standard image-based imitation learning policy [shafiullah2022behavior] from 300 demonstrations. These demos are generated using MimicGen [mandlekar2023mimicgen] with procedurally generated textures in 5 training room layouts. In each task, we use a Franka Panda robot mounted on an Omron mobile base initialized from the same base pose. Each task comes with a language description and a horizon of 500 steps. The goal of the benchmark is to deploy these policies into 5 held-out scenes with unseen room layouts and textures. Note that since these tasks come originally from RoboCasa, the tasks and policies are not designed for the problem of policy mobilization. Therefore, this benchmark can serve as an unbiased challenge task suite for testing methods for policy mobilization.
Quantifying mobilization feasibility of benchmark tasks. We plot the spatial mobilization feasibility metrics of the policies in the five simulated kitchen tasks in Fig. 2. We observe that the most challenging task to mobilize based on our metric is the Turn on Stove task ($\phi=0.031$). This reflects the difficulty to precisely turn a small knob which has limited contact area. Figure 3 shows the visual mobilization feasibility metric evaluated in the five environments. Results show that visual mobilization feasibilities do not necessarily match spatial mobilization feasibilities. More specifically, the Close Drawer task, which appears to be easy to mobilize spatially, turns out to be the most difficult in terms of visual mobilization feasibility metric. This is expected because even in spatially tolerant tasks, using proper visual features to find optimal policy initialization poses can be difficult. We show in experiments how the mobilization metrics are correlated with method performance in more detail.
3.3 Baseline Methods for Policy Mobilization
We include three baselines for policy mobilization in our Mobi-$\pi$ framework. We divide those into two categories. The first type of baselines navigates to the object of interest without considering the manipulation policy’s capabilities. These methods are not policy-aware. Representative methods include: (1) LeLaN [hiroselelan], a language-conditioned navigation model trained from large-scale videos. Since the pre-trained LeLaN model is trained on only real-world data, we fine-tune the pre-trained LeLaN model with 7.5k episodes of in-domain, randomized navigation data in simulation. At test time, we run LeLaN for a fixed number of steps to approach the object while using the scene point cloud for collision checking, and then execute the target policy. (2) VLFM [yokoyama2024vlfm], a zero-shot navigation model that explores the scene, uses a ground-truth detector to locate the object, selects a point on its unprojected point cloud, and navigates to it using RRT and the scene point cloud for collision checking. As shown in our experiments, these types of methods often fail to find suitable base poses as they are not taking manipulation policy performance into account, resulting in failed policy executions.
The second type of baselines is policy-aware and leverages large-scale data to connect navigation with manipulation. A representative method is BC w/ Nav: a Behavior Transformer [shafiullah2022behavior] trained to jointly perform navigation and manipulation using combined demonstrations. We provide 1,500 demonstrations per task, 5×5$\times$5 × the number of demos used compared to our base manipulation policy. We show in experiments that these methods generalize poorly to unseen layouts despite extensive data collection. See supplementary materials for baseline details.
@@ 4 A Novel Method for Policy Mobilization
We introduce a novel method for policy mobilization, illustrated in Figure 5. We want our method to satisfy the following criteria: (1) be policy-aware, (2) do not require extensive data collection for policy learning, and (3) operate in arbitrary unseen room layouts.
@@
4.1 Representing the Scene with 3D Gaussian Splatting
When the robot enters a previously unseen scene, it must first acquire a representation of the environment before identifying base poses from which the manipulation policy $\pi$ can be executed effectively. To this end, we employ 3D Gaussian Splatting (3DGS) [kerbl3Dgaussians], a technique that enables high-quality, novel-view synthesis from arbitrary viewpoints. Given a set of RGB-D images captured from various viewpoints, 3DGS constructs a scene representation composed of Gaussians $\mathcal{G}=\{g_{i}\}_{i=1}^{N_{g}}$ where each Gaussian $g_{i}=(\mu_{i},\Sigma_{i},o_{i},c_{i})$ encodes the 3D position $\mu_{i}\in\mathbb{R}^{3}$, spatial uncertainty via covariance matrix $\Sigma_{i}\in\mathbb{R}^{3\times 3}$, opacity $o_{i}\in\mathbb{R}$, and view-dependent color $c_{i}\in SH(4)$. To render an image from a novel camera pose $v\in SO(3)$, each Gaussian is projected to the image plane [zwicker2001ewa] and splatted using $\alpha$-blending to generate the final image $I=\text{GS}(v)$.
We construct the 3DGS model using 1,000 RGB-D images collected throughout the scene. This image acquisition process takes less than five minutes and involves driving the robot around the scene while varying camera height and viewing angle. We then reconstruct a point cloud from the depth images relative to the world frame to initialize the Gaussians. We then train the 3DGS models using nerfstudio [nerfstudio] for 30,000 update steps.
4.2 Hybrid Score Function for Estimating Policy Execution Viability
To estimate how likely a policy will be successful from a certain base pose $p$, we define a hybrid score function $K(p)$ that incorporates three key factors: (1) the rendered view should be in-distribution with respect to the training data, (2) the object of interest should be visible; (3) the robot pose should be free of collisions. To evaluate whether a rendered view is in-distribution, we leverage the initial camera observations $\{o_{1}^{k}\}_{k=1}^{N_{DS}}$ in the demonstration dataset. We compare the candidate view against these initial camera observations in a learned feature space. Among the image embeddings we evaluated, dense descriptors from DINO [caron2021emerging] proved most effective in distinguishing in- from out-of-distribution views and are robust to variations in scene texture. We precompute DINO features for all initial image observation, yielding$\{\text{DINO}(o_{1}^{k})\}_{k=1}^{N_{DS}}$. For a queried robot pose $p$, we render the corresponding image using our 3DGS models and compute its dense descriptors as $\text{DINO}(\text{GS}(f_{p\to v}(p)))$, where $f_{p\to v}$ denotes the transformation from base pose $p$ to camera pose $v$. The encoder $\text{DINO}(I)$ maps an image $I$ of size $H\times W$ to a feature tensor in $H^{\prime}\times W^{\prime}\times D$. To assess similarity, we apply K-nearest neighbors (KNN) in the feature space to compute the average L2 distance to the demonstration features, yielding an in-distribution score:
@@ 4.2 Hybrid Score Function for Estimating Policy Execution Viability
To increase robustness, we augment $K_{\text{id}}\in\{0,1\}$ with two additional binary checks. First, $K_{\text{id}}$ can return high scores for views that are texturally similar to the training set, even if they fail to include the task relevant object. To mitigate this, we verify object visibility using a vision-language model MiniCPM-v2 [yao2024minicpm] and define
@@
where $K_{\text{obj}}(p)=1$ if the object is detected and 00 otherwise. We found this approach to outperform conventional object detectors [minderer2022simple, liu2024grounding]. Second, we verify that pose $p$ is not in collision with the environment by consulting an occupancy map constructed from the depth data used the train the 3DGS model. We define:
@@
4.3 Robot Pose Optimization with Bayesian Optimization
@@
Given the score function $K(p)$, our objective is to find the robot base pose $p_{\pi}$ that maximizes the likelihood of successful policy execution: $p_{\pi}=\text{max}_{p}K(p)$. While some components of $K(p)$ are differentiable, we observe that the overall optimization landscape is non-smooth. In particular, small changes in the camera viewpoint can lead to abrupt variations in the score due to discontinuities introduced by occlusions, changes in object visibility or reachability. These factors make gradient-based optimization unreliable. To address this, we adopt a gradient-free optimization approach and use Bayesian Optimization (BO) to efficiently search for the optimal base pose $p_{\pi}$.

### v2

# Jingyun Yang†1, Isabella Huang∗2, Brandon Vu∗1 Max Bajracharya2, Rika Antonova3, Jeannette Bohg1

@@ Abstract
Learned visuomotor policies are capable of performing increasingly complex manipulation tasks. However, most of these policies are trained on data collected from limited robot positions and camera viewpoints. This leads to poor generalization to novel robot positions, which limits the use of these policies on mobile platforms, especially for precise tasks like pressing buttons or turning faucets. In this work, we formulate the policy mobilization problem: find a mobile robot base pose in a novel environment that is in distribution with respect to a manipulation policy trained on a limited set of camera viewpoints. Compared to retraining the policy itself to be more robust to unseen robot base pose initializations, policy mobilization decouples navigation from manipulation and thus does not require additional demonstrations. Crucially, this problem formulation complements existing efforts to improve manipulation policy robustness to novel viewpoints and remains compatible with them. We propose a novel approach for policy mobilization that bridges navigation and manipulation by optimizing the robot’s base pose to align with an in-distribution base pose for a learned policy. Our approach utilizes 3D Gaussian Splatting for novel view synthesis, a score function to evaluate pose suitability, and sampling-based optimization to identify optimal robot poses. To understand policy mobilization in more depth, we also introduce the framework, which includes: (1) metrics that quantify the difficulty of mobilizing a given policy, (2) a suite of simulated mobile manipulation tasks based on RoboCasa to evaluate policy mobilization, and (3) visualization tools for analysis. In both our developed simulation task suite and the real world, we show that our approach outperforms baselines, demonstrating its effectiveness for policy mobilization. Website: https://mobipi.github.io
@@ 1 Introduction
Recent advances in learning visuomotor manipulation policies have enabled robots to execute increasingly complex manipulation tasks using imitation and reinforcement learning [1, 2, 3, 4, 5, 6, 7]. However, most manipulation policies are trained on data collected from a limited set of robot base poses and camera viewpoints, leading to a tight coupling between the policy’s input distribution and the configurations seen during training. When such a policy is deployed on a mobile robot, changes in the robot’s base pose inherently alter its viewpoint. This can cause both the visual input and the reachable workspace to drift outside the policy’s training distribution, severely degrading performance.
Given the abundance of stationary robot data, it is highly desirable to reuse existing datasets and pre-trained policies for mobile manipulation. While retraining a policy to handle a wide range of base poses and viewpoints is a straightforward solution, it is also costly—requiring extensive data collection across diverse positions and scene variations [3, 8]. Moreover, as policies are deployed in mobile settings, their field of view expands, demanding stronger robustness to scene-level variations. An alternative strategy is to decouple navigation and manipulation by chaining a navigation module with the trained manipulation policy. However, most prior work in navigation specifies goals via language instructions or object locations [9, 10, 11, 12, 13, 14, 15, 16, 17, 18] rather than selecting base poses that are policy-compatible. As a result, the robot often navigates to viewpoints that lie outside of the policy’s training distribution, leading to unreliable performance.
@@
As a first attempt to tackle the policy mobilization problem, we propose a method that searches for robot base poses likely to yield successful policy execution. Our approach (a) represents the scene using 3D Gaussian Splatting [19], (b) uses differentiable rendering to evaluate whether a candidate pose is in-distribution, provides visibility of task-relevant objects, and avoids collisions, and (c) optimizes over poses with a sampling-based procedure to identify the most suitable robot base location. To systematically study the policy mobilization problem, we present the ramework, which includes (1) metrics to quantify the difficulty of mobilizing a given policy, (2) a suite of simulated mobile manipulation tasks built on RoboCasa [20], (3) visualization tools for qualitative analysis.
We first evaluate non-policy aware baselines that decouple navigation and manipulation without considering that policies perform better from specific base poses. These approaches frequently fail due to the poor alignment with the policy’s training distribution. We also examine policy-aware baselines that learn to connect navigation and manipulation using large amounts of additional training data and show that they still struggle to generalize to unseen room layouts. Finally, we validate our approach on both the imulation suite and three real-world mobile manipulation tasks. Our method outperforms all baselines in various simulation and real-world experiments, enabling effective deployment of manipulation policies trained solely on stationary robot data.
@@ 2 Related Work
Semantic navigation. In semantic navigation, robots navigate environments to find specific objects or regions. Prior work [21, 22, 23, 9, 24, 10, 25] uses classical methods like SLAM [26] to build and explore semantic maps, while recent methods use learning-based approaches [11, 12, 13, 14, 16, 17]. Other methods predict semantic top-down maps [27, 28, 18] and navigate using waypoint planners [29]. These approaches typically stop in an area where the object is visible and close, without ensuring that the viewpoint is suitable for manipulation. In contrast, our method selects navigation targets that maximize downstream manipulation policy success.
@@
Imitation learning. Imitation learning teaches robots tasks from demonstrations [30, 31, 32, 1, 33, 34, 35], with recent efforts scaling to generalist models [3, 4, 36, 37, 5, 8, 6, 38, 7]. Training such policies for both navigation and manipulation requires massive data. Instead, we use imitation learning for the manipulation policy and leverage existing navigation methods to move the robot to an in-distribution robot pose ensuring policy success. Our insight is that by selecting suitable poses for manipulation, we reduce the need for collecting more data from many robot poses or viewpoints and improve deployment in novel scenes. We focus on single-task policies and leave multi-task extensions to future work.
@@
Mobile manipulation. Prior work has studied various ways to bring learned policies to mobile manipulation [39, 40, 41, 42, 34, 43, 44, 45]. Some methods [3, 34] rely on end-to-end learning but require large-scale data. Others [39, 40, 41, 43] assume known initial base poses for a given task. ARPlace [46] predicts valid robot poses for running skills but focuses more on reachability rather than viewpoint restrictions for these skills. ASC [47] trains coordinated mobile pick-and-place skills with RL in simulation but requires highly accurate simulators, which are often difficult to obtain for complex real-world tasks with deformables or complex multi-object interactions. Some works use reachability maps or abstract planners [48, 49], but ignore that policies may be viewpoint sensitive. Our approach directly addresses this by finding robot base and camera poses that enable manipulation policy success. This “mobilizes” policies with fixed initial base poses, making them usable on mobile robots in complex environments, including those with deformable objects or multi-object interactions.
@@
In-distribution detection. Policy mobilization involves finding an in-distribution initial condition for a given policy. Prior work has extensively studied in-distribution detection or anomaly detection, covering classification-based methods [50, 51, 52, 53, 54], reconstruction methods [55, 56], methods that utilize generative adversarial networks (GANs) [57, 58], and diffusion-based methods [59]. Nearly all these methods assume a passive setting where they detect whether a given input is in-distribution without solving the problem of how an agent can find a sample that is more in-distribution. Recent robotics works have started to address out-of-distribution (OOD) detection [60, 61, 62] but treat the perception module as a passive component without actively updating the robot’s camera or base pose to mitigate OOD scenarios. Our work closes this gap, marrying in-distribution detection with active robot pose optimization in a mobile manipulation context.
@@
Hierarchical methods for policy learning. Our setting resembles Hierarchical Reinforcement Learning (HRL), where an option [63] includes an initiation set, a policy, and a termination function. HRL methods [64, 65, 66, 67, 68] often learn these jointly but have limited success in complex settings. Instead of learning all elements of an option, we focus on selecting optimal robot poses for pre-trained manipulation policies in realistic environments.
@@
Training view-robust policies. Related to the problem setting of our work, prior work has studied methods of obtaining view-robust manipulation policies. One approach achieves this by collecting multi-view data in simulation [69, 70, 71], but suffers from the sim-to-real transfer challenge. Some works achieve view-robustness by using 3D representations, but training a policy to use 3D input requires data with well-calibrated camera extrinsics [72, 73, 74], which is not readily available in many robot learning datasets today. Some works achieve real-world view robustness by manually collecting multi-view data in the real world [8], but this requires extensive human effort. There are proposals for using single-view novel view synthesis to bypass the need for multi-view data [75], but methods for novel view synthesis from a single image struggle to effectively render views that are too different from the given viewpoint. From a view robustness perspective, our work takes advantage of the mobility of the robot so the robot can navigate to a base pose with an appropriate camera viewpoint for subsequent policy execution, thus requiring no extensive additional data collection.
3 The Policy Mobilization Problem
Let the environment be an MDP $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R})$. $\mathcal{S}$ is the state space, where a state $s=(s^{\text{env}},p,q)$ consists of the environment state $s^{\text{env}}$, the mobile base pose $p\in SE(2)$, and the arm configuration $q\in\mathbb{R}^{n_{\text{dof}}}$ with $n_{\text{dof}}$ denoting the number of degrees of freedom of the robot arms. $\mathcal{A}$ is an action space that consists of base and arm actions. $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,+\infty)$ represents the state transition probability. $\mathcal{R}:\mathcal{S}\rightarrow\mathbb{R}$ is the reward function. We are given a pre-trained manipulation policy $\pi:\mathcal{O}\rightarrow\mathcal{A}$, where $\mathcal{O}$ is the observation space. We measure performance of the mobilized policy with the expected final reward $J(\pi|s_{\text{start}})=\mathbb{E}_{\tau\sim\pi}\big[\mathcal{R}(s_{H})|s_{0}=s_{\text{start}}\big]$, where $\tau=(s_{0},a_{0},s_{1},\cdots,s_{H})$ is a trajectory of length $H$ generated by executing $\pi$ from state $s_{\text{start}}$. The initial state is $s_{0}=(s^{\text{env}}_{0},p_{0},q_{0})$. Policy mobilization aims to find an optimal base pose $p^{*}$ to run the manipulation policy such that the expected return is maximized: $p^{*}=\text{argmax}_{p\in\text{SE}(2)}J\big(\pi|(s^{\text{env}}_{0},p,q_{0})\big).$
@@ 4 A Novel Method for Policy Mobilization
We introduce a novel method for policy mobilization, illustrated in Figure 2. We want our method to satisfy the following criteria: (1) take into account the capabilities of the policy, (2) do not require extensive data collection for policy learning, and (3) operate in arbitrary unseen room layouts.
@@
Representing the scene with 3D Gaussian Splatting. When the robot enters a previously unseen scene, it must first acquire a representation of the environment before identifying base poses from which the manipulation policy $\pi$ can be executed effectively. To this end, we employ 3D Gaussian Splatting (3DGS) [19], a technique that enables high-quality, novel-view synthesis from arbitrary viewpoints. Given a set of RGB-D images captured from various viewpoints, 3DGS constructs a scene representation composed of Gaussians $\mathcal{G}=\{g_{i}\}_{i=1}^{N_{g}}$ where each Gaussian $g_{i}=(\mu_{i},\Sigma_{i},o_{i},c_{i})$ encodes the 3D position $\mu_{i}\in\mathbb{R}^{3}$, spatial uncertainty via covariance matrix $\Sigma_{i}\in\mathbb{R}^{3\times 3}$, opacity $o_{i}\in\mathbb{R}$, and view-dependent color $c_{i}\in SH(4)$. To render an image from a novel camera pose $v\in SO(3)$, each Gaussian is projected to the image plane [76] and splatted using $\alpha$-blending to generate the final image $I=\text{GS}(v)$.
We construct the 3DGS model using 1,000 RGB-D images collected throughout the scene. This image acquisition process takes less than 5 minutes and involves driving the robot around the scene while varying the camera height and viewing angle. Future work may consider automating this process so that no human is involved for this image collection. We then reconstruct a point cloud from the depth images relative to the world frame to initialize the Gaussians. Finally, we train the 3DGS models using nerfstudio [77] for 30,000 update steps.
[…]
@@
To increase robustness, we augment $K_{\text{id}}\in\{0,1\}$ with two additional binary checks. First, $K_{\text{id}}$ can return high scores for views that are texturally similar to the training set, even if they fail to include the task relevant object. To mitigate this, we verify object visibility using a vision-language model MiniCPM-v2 [79] and define
@@
where $K_{\text{obj}}(p)=1$ if the object is detected and 0 otherwise. We found this approach to outperform conventional object detectors [80, 81]. Second, we verify that pose $p$ is not in collision with the environment by consulting an occupancy map constructed from the depth data used to train the 3DGS model. We define:
@@
Robot pose optimization with Bayesian Optimization. Given the score function $K(p)$, our objective is to find the robot base pose $p_{\pi}$ that maximizes the likelihood of successful policy execution: $p_{\pi}=\text{max}_{p}K(p)$. While some components of $K(p)$ are differentiable, we observe that the overall optimization landscape is non-smooth. In particular, small changes in the camera viewpoint can lead to abrupt variations in the score due to discontinuities introduced by occlusions, changes in object visibility or reachability. These factors make gradient-based optimization unreliable. To address this, we adopt a gradient-free optimization approach and use Bayesian Optimization (BO) to efficiently search for the optimal base pose $p_{\pi}$.